# HuB: Learning Extreme Humanoid Balance

### v1

@@ 3.3 Balance-Aware Policy Learning
Balance Shaping Rewards. Merely relaxing the tracking constraint does not guarantee that the policy will learn physically feasible behaviors. Without structured guidance, reinforcement learning can converge to suboptimal solutions that violate task intent or physical plausibility. To handle this challenge, we design a set of shaping rewards. (i) Center of mass (COM) reward encourages the horizontal projection of the COM to remain within the support polygon, thereby helping the policy learn to adjust its pose to satisfy balance constraints. (ii) Foot contact mismatch penalty categorizes the landing state of each foot as either in contact with the ground or not, and penalizes discrepancies between the humanoid’s and the reference’s contact states. For example, it discourages the non-supporting foot from making unintended ground contact during single-leg balancing. While such contact may offer momentary stability, it violates task constraints and compromises the integrity of the intended balance behavior. (iii) Close feet penalty prevents the feet from getting too close to one another, reducing the risk of inter-foot collisions and encouraging more stable lower-body poses. Collectively, these shaping rewards promote the emergence of balance-aware, physically plausible motion policies. More reward details are described in Section B.2.
@@ 4.2 Simulation Results
Retargeting Results. To assess the impact of SMPL-initialization on optimization-based retargeting, we compare the retargeting loss after 500 optimization steps between solutions optimized with and without SMPL-initialization. As shown in the Figure 3, solutions with SMPL-initialization consistently achieve lower losses across all tasks, with notably large reductions in Deep Squat.
@@ Acknowledgments
We thank Jiacheng You and Haoyang Weng for their helpful discussions, and Junming Zhao and Sicong Dai for their assistance with the real-world experiments.

### v2

@@ 3.3 Balance-Aware Policy Learning
Balance Shaping Rewards. Merely relaxing the tracking constraint does not guarantee that the policy will learn physically feasible behaviors. Without structured guidance, reinforcement learning can converge to suboptimal solutions that violate task intent or physical plausibility. To handle this challenge, we design a set of shaping rewards. (i) Center of mass (COM) reward encourages the vertical projection of the COM to remain within the support polygon, thereby helping the policy learn to adjust its pose to satisfy balance constraints. (ii) Foot contact mismatch penalty categorizes the landing state of each foot as either in contact with the ground or not, and penalizes discrepancies between the humanoid’s and the reference’s contact states. For example, it discourages the non-supporting foot from making unintended ground contact during single-leg balancing. While such contact may offer momentary stability, it violates task constraints and compromises the integrity of the intended balance behavior. (iii) Close feet penalty prevents the feet from getting too close to one another, reducing the risk of inter-foot collisions and encouraging more stable lower-body poses. Collectively, these shaping rewards promote the emergence of balance-aware, physically plausible motion policies. More reward details are described in Section B.2.
@@ 4.2 Simulation Results
Retargeting Results. To assess the impact of SMPL-initialization on retargeting, we compare the retargeting loss after 500 optimization steps between solutions optimized with and without it. As shown in Figure 3, solutions with SMPL-initialization consistently achieve lower losses across all tasks, with notably large reductions in Deep Squat.
@@ Acknowledgments
We thank Jiacheng You, Haoyang Weng, and Bike Zhang for their helpful discussions, and Junming Zhao and Sicong Dai for their assistance with the real-world experiments. This work is supported by the National Key R&D Program of China (2022ZD0161700), National Natural Science Foundation of China (62176135, 12201341), Shanghai Qi Zhi Institute Innovation Program SQZ202306 and the Tsinghua University Dushi Program.
@@ B.5 Hyperparameters
B.6 Foot Contact Labeling
In the grounded foot correction stage, the foot with the lower ankle height is considered the grounded foot. For the foot contact mismatch penalty, please refer to Table 5 for the criteria used to determine which foot is considered grounded.
@@ C.2 Additional Results
To validate the necessity of our IMU-Centric Observation Perturbation component, we conduct simulation experiments on Bruce Lee’s Kick comparing our coupled OU noise with other variants, including coupled Uniform noise as well as independent OU and independent Uniform (vanilla) noise. Here, coupled means noise is first applied to the IMU’s orientation observation and then the policy input observation is computed based on this noisy orientation, while independent means noise is directly added to the policy input observation. The results in Table 9 show the coupled OU noise achieves the best overall performance, supporting its ability to better model real sensor noise and enhance policy robustness.
| Completion | Stability | Tracking Error | ||||||
| Method | ||||||||
| Independent Uniform | 98 | 0.02 | 162.63 | 6.10 | 4.40 | 105.21 | 4.81 | 3.12 |
| Independent OU | 98 | 0.03 | 140.68 | 5.34 | 3.35 | 99.22 | 4.55 | 2.82 |
| Coupled Uniform | 100 | 0.00 | 80.02 | 1.57 | 0.50 | 69.45 | 3.45 | 2.31 |
| Coupled OU (ours) | 100 | 0.00 | 76.44 | 1.28 | 0.46 | 67.18 | 3.31 | 2.33 |
C.3 Failure Analysis of Baselines
In our experiments, the failures of Baselines (H2O and OmniH2O) are primarily due to either falling or foot contact mismatch. The latter is treated as a failure, as contact with both feet indicates the single-leg task is not successfully completed. Refer to the Comparative Results section on the project website for videos: the robot fails the Swallow Balance due to falling, and fails the Ne Zha Pose and Single-Leg Stand due to contact mismatch. The low success rates of the baselines can be attributed to many factors, including—but not limited to—the absence of a contact penalty, overly small tracking tolerances, and inappropriate reward designs such as penalizing non-flat base orientations.